If Kelly's budget has an X intercept of 40 and a Y intercept of 60, and PY is $20,
a. What is Kellys income?
b. What is Price of good X?
c. What is the slope of the budget line?
1
Expert's answer
2017-02-21T13:51:10-0500
a. X*Px+Y*Py=I
0*Px+60*20=I
I=120

b. 40*Px+0*20=120
Px=30

c. Slope=-Px/Py=- 30/20=-1.5